OpenStudy (pphalke):

Randomly drawn from a card deck of 52 cards. Find the probability of drawing a face card i still dont understand this

OpenStudy (pphalke):

@mayankdevnani @habbababba1

OpenStudy (pphalke):

@Kayla.M23

OpenStudy (welshfella):

how many face cards are there in the pack of 52?

OpenStudy (habbababba1):

You are actually asking about a conditional probability of an event given the other. A standard conditional probability answer: Randomly drawing one card, each of the 54 cards in a deck has the same probability of being drawn, i.e., 1 54 154 . Since there are 13 diamond cards, the probability of drawing a diamond card is 1 54 ∗ 13 154∗13 ; there are 3 cards that are both face and diamond, so the probability of drawing such a card is 3/54. If you use the conditional probability formula for event B (drawing a card that is diamond and face) given event A (drawing a card that is a diamond), P ( B | A ) = P ( B ∩ A ) P ( A ) = 3 / 54 13 / 54 = 3 / 13 P(B|A)=P(B∩A)P(A)=3/5413/54=3/13 . A shortcut way to solve this: since the drawn card is constraint to be a diamond, each of the 13 diamond cards have equal chances of being the drawn card; as only 3 of them are face cards, the probably of actually getting a face card is simply 3/13.

OpenStudy (habbababba1):

i can count cards so this is a perfect question for me haha

OpenStudy (welshfella):

hmmm I can't see how you make this out to be conditional probability

OpenStudy (pphalke):

@habbababba1 is that the answer

OpenStudy (habbababba1):

what are your options

OpenStudy (habbababba1):

and yes that is the answer

OpenStudy (habbababba1):

3/13